OpenStudy (anonymous):

What is the midpoint of (-2, 0) and (4, -2)?

OpenStudy (ivettef365):

use midpoint formula

hartnn (hartnn):

The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

Thank you, couldnt find a website to help with it.

hartnn (hartnn):

\(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

its (1,-1)

OpenStudy (anonymous):

Thank you!